2. Imagine a closed economy where the dictatorship does not allow any trade with other countries. If the autonomous consumption equals R10 mil., autonomous investment equals R20 mil., the tax rate equals 15%, government expenditure R30 mil. and the marginal propensity to consume (MPC = c) equals 0.7.


2.1 What is the size of the multiplier ()?


2.2 What is the equilibrium level of income (Y0)?


2.3 How much will the tax revenue for the goverment be in this economy?


2.4 How much will the disposable income be in this economy?



1
Expert's answer
2021-11-02T10:21:37-0400

2.1 The size of the multiplier is:

"m = \frac{1}{1 - 0.7(1 - 0.15)} = 2.47."

2.2 The equilibrium level of income (Y0) is:

"Y0 = 10 + 0.7(1 - 0.15)Y0 + 30 + 20 = 60 + 0.595Y0,"

0.405Y0 = 60,

Y0 = 148.15.

2.3 How much will the tax revenue for the goverment be in this economy?

"T = 0.15Y0 = 0.15*148.5 = 22.22."

2.4 The disposable income will be:

"DY = Y0 - T = 148.15 - 22.22 = 125.93."